OpenStudy (anonymous):

what is the equation for -4(2n+3)=24

OpenStudy (anonymous):

you want to solve for n?

OpenStudy (anonymous):

-4n-12=24 -4n=36 n=-9

OpenStudy (anonymous):

Sorry, that is incorrect

OpenStudy (anonymous):

-8n-12=24 -8n=36 n=-4.5

OpenStudy (anonymous):

Didn't have my glasses :(

OpenStudy (anonymous):

Thanks hedyeh99

OpenStudy (anonymous):

Your welcome :D